# Hierarchical Deep Learning for Diatom Image Classification: A Multi-Level Taxonomic Approach

### v1

@@ Abstract
Accurate taxonomic identification of diatoms is essential for aquatic ecosystem monitoring, yet conventional methods depend heavily on expert taxonomists. Recent deep learning approaches improve automation, but most treat diatom recognition as flat classification predicting only one taxonomic rank. We investigate whether embedding taxonomic hierarchy into neural network architectures can improve both accuracy and error locality.
We introduce a hierarchical convolutional network with five cascaded heads that jointly predict class, order, family, genus, and species. Each head receives shared backbone features and probability distributions from higher levels, with binary masks restricting predictions to valid descendants during training and inference. Using a filtered dataset of 1,456 diatom images covering 82 species, we compare hierarchical and flat models under identical settings.
The hierarchical model matches flat baselines at species level (69.4% accuracy) while outperforming at all upper taxonomic levels. When species predictions fail, errors remain taxonomically local: 92.5% of misclassified species are correctly predicted at genus level, versus 67.2% for flat baselines. The hierarchical model reduces mean taxonomic distance by 38.2% (1.209 vs. 1.955).
@@ I-A Background and Motivation
Diatoms are unicellular microalgae with ornate silica cell walls, and shifts in their communities provide sensitive indicators of environmental change [round1990, kelly1995]. Reliable taxonomic identification is therefore central to water-quality assessment, ecological monitoring, and biodiversity studies [krammer2012, stevenson1999]. Yet the standard workflow still depends on expert taxonomists who inspect microscope images by hand, a process that is time-consuming, labor-intensive, and prone to inter-observer variability [wang2014diatom]. As monitoring programs continue to scale up, demand for automated and trustworthy identification pipelines keeps growing.
Recent progress in deep learning has improved automated diatom classification [krizhevsky2012imagenet, he2016deep, simonyan2015very]. Convolutional neural networks (CNNs) with transfer learning reach high accuracy at individual taxonomic ranks [gonzalez2025, scdiatomnet2024, deng2009imagenet, pan2010survey], learning discriminative visual features directly from microscope imagery and removing the need for hand-crafted descriptors. Almost all of these approaches, however, treat the task as flat classification: the network predicts a single rank, typically genus or species, and the upper-level labels are obtained by traversing the taxonomy tree. This setup is simple and efficient, but it leaves open a central question: can architectures that embed the taxonomic hierarchy deliver tangible benefits over single-level prediction with post-hoc tree lookup?
@@ I-B Research Gap and Objectives
    Hierarchical models: Multi-output CNNs with five cascaded heads (class, order, family, genus, and species). Each head receives the backbone features concatenated with the probability distributions from all higher levels, implementing feature fusion between visual and hierarchical cues. During training we use teacher forcing with ground-truth masks, which creates bidirectional information flow: (1) bottom-up through gradient backpropagation as all five losses update the shared backbone, and (2) top-down through hierarchical masks that constrain each level’s output space. During inference we switch to greedy hierarchical prediction: the argmax at each level determines the mask for the next level.
@@ II-A Automated Diatom Classification
Automated diatom identification has evolved significantly over the past two decades. Early approaches relied on hand-crafted features such as shape descriptors, texture analysis, and morphological measurements coupled with classical classifiers including k-nearest neighbors, support vector machines, and decision trees [pedraza2017, dimitrovski2012, chaushevska2020]. These methods encoded valve outline, striae density, and ornamentation into engineered feature vectors, which required substantial domain expertise and were sensitive to variations in focus, illumination, and specimen damage. Performance remained limited on large, taxonomically diverse datasets and did not scale well to hundreds of species.
The advent of deep learning has transformed diatom image analysis. Convolutional neural networks (CNNs), particularly in combination with transfer learning from large natural-image datasets such as ImageNet [deng2009imagenet, krizhevsky2012imagenet], have yielded substantial accuracy improvements over hand-crafted pipelines on both genus and species level tasks [gonzalez2025, scdiatomnet2024, pan2010survey]. Modern CNN architectures such as VGGNet, Inception, ResNet, and EfficientNet learn hierarchical representations automatically through successive convolutional and pooling layers [simonyan2015very, szegedy2015going, he2016deep, tan2019efficientnet, zeiler2014visualizing]. Transfer learning leverages features learned on large-scale datasets to improve performance on specialized domains with limited training data [yosinski2014transferable, pan2010survey, tan2018survey].
@@ II-B Hierarchical Classification Methods
Multi-task learning [caruana1997multitask, goodfellow2016], where models jointly optimize multiple related objectives, provides a theoretical foundation for hierarchical architectures that predict all taxonomic levels simultaneously. When multiple related tasks share representations through a common backbone, joint optimization can improve individual task performance through complementary learning signals. Recent work in biological classification has demonstrated the effectiveness of shared-backbone architectures for multi-level taxonomic prediction [elhamod2021hierarchy, wittich2018].
@@ II-C Hierarchical Classification in Biology
Beyond diatoms, hierarchical deep models have been investigated for plant and leaf identification. Wittich [wittich2018] trained CNNs to predict plant family and genus independently, demonstrating that models can generalize taxonomic knowledge from seen to unseen species. Elhamod et al. [elhamod2021hierarchy] proposed Hierarchy-Guided Neural Networks (HGNN), which incorporate genus-level information into species-level feature learning through parallel prediction branches. Studies in this area report that enforcing taxonomic structure (either via multi-task learning across ranks or through constrained decoding) reduces implausible errors and improves the usability of automated tools for field biologists [wittich2018, elhamod2021hierarchy].
Nevertheless, to the best of our knowledge, end-to-end CNN architectures that embed taxonomic constraints at every prediction level have not yet been systematically evaluated for diatom classification. Existing deep-learning-based diatom systems remain predominantly flat [gonzalez2025, scdiatomnet2024], while existing hierarchical diatom methods rely on hand-crafted features or non-differentiable ensembles [dimitrovski2012, chaushevska2020], leaving a gap between flat CNN architectures and end-to-end hierarchical models with explicit structural constraints in this domain.
III Materials and Methods
III-A Dataset and Data Pipeline
III-A1 Data Cleaning and Taxonomy Construction
The initial raw dataset contains 4,881 microscope images captured from System for the Identification of European Diatoms [krammer2012], with taxonomic labels manually annotated by research staff and students at the Yantai Institute of Coastal Zone Research, Chinese Academy of Sciences, under the supervision of Dr. Yin-chu Wang. The dataset spans 4 classes, 15 orders, 28 families, 47 genera, and 684 species, annotated following Round et al. [round1990]. We apply systematic data cleaning to correct taxonomic inconsistencies: (1) merge the rare class Mediophyceae into Coscinodiscophyceae [round1990], (2) remove samples with incomplete annotations at class, order, or family levels, (3) exclude species with uncertain labels containing markers such as “sp.”, “cf.”, or “?”, and (4) resolve hierarchical conflicts where taxa map to multiple parents using majority voting. After cleaning, the dataset contains 3,522 samples across 3 classes, 13 orders, 23 families, 36 genera, and 608 species.
III-A2 Progressive Minimum Sample Filtering
To support progressive training and fair comparison, we apply a uniform threshold of 10 images per taxon at each taxonomic level, yielding four nested datasets: H-CO (3,506 samples, filters order), H-COF (3,495 samples, filters order and family), H-COFG (3,472 samples, filters order, family, and genus), and H-COFGS (1,456 samples, filters all five levels). The final filtered dataset H-COFGS, used by both hierarchical models and the flat baseline F-S, spans 3 classes, 9 orders, 13 families, 17 genera, and 82 species.
All datasets are randomly split into training (70%), validation (15%), and test (15%) sets, stratified by species with a fixed random seed (42) to ensure reproducibility. Input images are resized to $320\times 320$ pixels and normalized using ImageNet statistics. Data augmentation during training includes random horizontal flips, rotations within $\pm 15^{\circ}$, and color jittering [shorten2019survey, krizhevsky2012imagenet].
Figure 2: Hierarchical diatom classification architecture (H-COFGS). The EfficientNet-B0 backbone extracts features from input images, which are fed to five cascaded classification heads (Class, Order, Family, Genus, Species). Each head receives backbone features concatenated with probability distributions from all parent levels. During training (left), ground-truth labels are used for masking (teacher forcing); during inference (right), predicted labels determine masks at each level (greedy hierarchical prediction). Input diatom image adapted from [berezovska2015diatom], $C$ BY-SA 4.0.
@@ III-B2 Hierarchical Models
Hierarchical models explicitly embed taxonomic relationships by predicting labels at multiple levels in a single forward pass through cascaded classification heads (Fig. 2). Each head at level $\ell$ receives concatenated features from the backbone and all parent-level probability distributions:
$\mathbf{x}_{\ell}=[\mathbf{f}_{\text{backbone}};\mathbf{p}_{\text{class}};\mathbf{p}_{\text{order}};\ldots;\mathbf{p}_{\ell-1}],$ (1)
implementing feature fusion between visual and hierarchical cues. As illustrated in Fig. 2, head architectures scale with task complexity: the class head uses a two-layer MLP, intermediate levels (order, family) use three-layer MLPs, and fine-grained levels (genus, species) use four-layer architectures (detailed in Table I).
A key design is the masking mechanism that enforces taxonomic consistency. At each level $\ell$, the model produces raw logits $\mathbf{z}_{\ell}$ transformed via masked softmax with a binary mask $\mathbf{m}_{\ell}\in\{0,1\}^{n_{\ell}}$:
@@
where $\mathbf{m}_{\ell}$ is constructed by querying the taxonomy tree: only child classes at level $\ell$ that belong to parent class $c_{\ell-1}$ receive a mask value of 1.
For hierarchical models, each taxonomic level receives a separate focal loss [lin2017focal], and the total loss is a weighted sum:
@@
where weights increase progressively from coarse to fine levels (class: 0.8, order: 0.9, family: 1.0, genus: 1.2, species: 1.5) to emphasize harder, fine-grained predictions. Focal loss with $\alpha=0.25$ and $\gamma=2.0$ addresses class imbalance [lin2017focal, buda2018systematic].
III-C Training Strategy
III-C1 Training and Validation
All models are trained using the AdamW optimizer [loshchilov2017decoupled] with an initial learning rate of $5\times 10^{-4}$, weight decay $1\times 10^{-4}$, a ReduceLROnPlateau scheduler (decay factor 0.5, patience 5 epochs), batch size 32, a maximum of 80 epochs, and early stopping with a patience of 15 epochs based on validation performance (deepest-level weighted F1 score).
Standard deep learning training typically follows four steps: (1) a forward pass that maps inputs to logits, (2) loss computation based on logits and ground-truth labels, (3) a backward pass that propagates gradients through the network, and (4) parameter updates performed by an optimizer. In our hierarchical setting, the forward pass is specialized: logits are computed sequentially from coarse to fine levels, and each head receives a feature vector obtained by concatenating the shared backbone representation with all previous levels’ probability distributions (Equation 1), so that hierarchical information flows explicitly from parent to child. During loss computation, we further exploit the taxonomy by applying ground-truth-derived hierarchical masks to the logits at each level and then feeding these masked logits into a weighted focal loss (Equation 3), effectively restricting every level’s candidate space to valid children of its parent and stabilizing training. The subsequent backward pass aggregates gradients from all levels and propagates them through the shared backbone, yielding bidirectional information flow: bottom-up via gradient backpropagation and top-down via the hierarchical constraints imposed during loss computation.
@@ III-C1 Training and Validation
We adopt a progressive training strategy across taxonomic depths. First, the flat class-only baseline F-C is trained from ImageNet initialization on the full cleaned dataset (3522 samples). Next, H-CO initializes its backbone from F-C and is trained on the class–order dataset H-CO (labels_clean_CO.csv, 3506 samples), which is filtered to satisfy the minimum-sample threshold at the order level. Similarly, H-COF initializes its backbone from H-CO and trains on the H-COF dataset (labels_clean_COF.csv, 3495 samples), H-COFG initializes from H-COF and trains on H-COFG (labels_clean_COFG.csv, 3472 samples), and H-COFGS initializes from H-COFG and trains on H-COFGS (labels_clean_COFGS.csv, 1456 samples), each using its own filtered dataset that enforces the minimum-sample threshold at its deepest taxonomic level.
III-C2 Inference
During test inference, flat models predict a single taxonomic level directly, with labels at other levels obtained by querying the taxonomy tree. Hierarchical models use the same greedy hierarchical prediction as validation: logits are computed sequentially (Equation 1), probability distributions are computed via softmax, and we greedily select the class with the highest probability (argmax) at each level, using the predicted index from the previous level to mask the current level’s probability distribution via Equation 2, ensuring biological consistency:

### v2

@@ Abstract
Accurate taxonomic identification of diatoms is essential for aquatic ecosystem monitoring, yet conventional methods depend heavily on expert taxonomists. Recent deep learning approaches improve automation, but most treat diatom recognition as flat classification, predicting only one taxonomic rank. We investigate whether embedding taxonomic hierarchy into neural network architectures can improve both accuracy and error locality.
We introduce DiatomCascadeNet (H-COFGS), a hierarchical convolutional network with five cascaded heads that jointly predict class, order, family, genus, and species. Each head receives shared backbone features and probability distributions from higher levels, with binary masks restricting predictions to valid descendants during training and inference. Using a filtered dataset of 1,456 diatom images covering 82 species, we compare hierarchical and flat models under identical settings.
H-COFGS matches flat baselines at the species level (69.4% accuracy) while outperforming at all upper taxonomic levels. When species predictions fail, errors remain taxonomically local: 92.5% of misclassified species are correctly predicted at the genus level, versus 67.2% for flat baselines. H-COFGS reduces mean taxonomic distance by 38.2% (1.209 vs. 1.955).
@@ I-A Background and Motivation
Diatoms are unicellular microalgae with ornate silica cell walls, and shifts in their communities provide sensitive indicators of environmental change [round1990, kelly1995]. Reliable taxonomic identification is therefore central to water-quality assessment, ecological monitoring, and biodiversity studies [krammer2012, stevenson1999]. Yet the standard workflow still depends on expert taxonomists who inspect microscope images by hand, a process that is time-consuming, labor-intensive, and prone to inter-observer variability. As monitoring programs continue to scale up, demand for automated and trustworthy identification pipelines keeps growing.
Recent progress in deep learning has improved automated diatom classification [krizhevsky2012imagenet, he2016deep, simonyan2015very]. Convolutional neural networks (CNNs) with transfer learning from ImageNet [deng2009imagenet] reach high accuracy at individual taxonomic ranks [gonzalez2025, diatomnet2024], learning discriminative visual features directly from microscope imagery and removing the need for hand-crafted descriptors [pan2010survey]. Almost all of these approaches, however, treat the task as flat classification: the network predicts a single rank, typically genus or species, and the upper-level labels are obtained by traversing the taxonomy tree. This setup is simple and efficient, but it leaves open a central question: can architectures that embed the taxonomic hierarchy deliver tangible benefits over single-level prediction with post-hoc tree lookup?
@@ I-B Research Gap and Objectives
    Hierarchical models: Multi-output CNNs with five cascaded heads (class, order, family, genus, and species). Each head receives the backbone features concatenated with the probability distributions from all higher levels, implementing feature fusion between visual and hierarchical cues. During training, we use teacher forcing with ground-truth masks, which creates bidirectional information flow: (1) bottom-up through gradient backpropagation as all five losses update the shared backbone, and (2) top-down through hierarchical masks that constrain each level’s output space. During inference, we switch to greedy hierarchical prediction: the argmax at each level determines the mask for the next level.
@@ II-A Automated Diatom Classification
Automated diatom identification has evolved significantly over the past two decades. Early approaches relied on hand-crafted features such as shape descriptors, texture analysis, and morphological measurements coupled with classical classifiers, including k-nearest neighbors, support vector machines, and decision trees [pedraza2017, dimitrovski2012, chaushevska2020]. These methods encoded valve outline, striae density, and ornamentation into engineered feature vectors, which required substantial domain expertise and were sensitive to variations in focus, illumination, and specimen damage. Performance remained limited on large, taxonomically diverse datasets and did not scale well to hundreds of species.
The advent of deep learning has transformed diatom image analysis. Convolutional neural networks (CNNs), particularly in combination with transfer learning from large natural-image datasets such as ImageNet [deng2009imagenet, krizhevsky2012imagenet], have yielded substantial accuracy improvements over hand-crafted pipelines on both genus and species level tasks [gonzalez2025, diatomnet2024]. Modern CNN architectures such as VGGNet, Inception, ResNet, and EfficientNet learn hierarchical representations automatically through successive convolutional and pooling layers [simonyan2015very, szegedy2015going, he2016deep, tan2019efficientnet, zeiler2014visualizing]. Transfer learning leverages features learned on large-scale datasets to improve performance on specialized domains with limited training data [yosinski2014transferable, pan2010survey, tan2018survey].
@@ II-B Hierarchical Classification Methods
Multi-task learning [caruana1997multitask, goodfellow2016], where models jointly optimize multiple related objectives, provides a theoretical foundation for hierarchical architectures that predict all taxonomic levels simultaneously. When multiple related tasks share representations through a common backbone, joint optimization can improve individual task performance through complementary learning signals. Recent work in biological classification has demonstrated the effectiveness of shared-backbone architectures for multi-level taxonomic prediction [elhamod2022hierarchy, wittich2018].
@@ II-C Hierarchical Classification in Biology
Beyond diatoms, hierarchical deep models have been investigated for plant and leaf identification. Wittich [wittich2018] trained CNNs to predict plant family and genus independently, demonstrating that models can generalize taxonomic knowledge from seen to unseen species. Elhamod et al. [elhamod2022hierarchy] proposed Hierarchy-Guided Neural Networks (HGNN), which incorporate genus-level information into species-level feature learning through parallel prediction branches. Studies in this area report that enforcing taxonomic structure (either via multi-task learning across ranks or through constrained decoding) reduces implausible errors and improves the usability of automated tools for field biologists [wittich2018, elhamod2022hierarchy].
Nevertheless, to the best of our knowledge, end-to-end CNN architectures that embed taxonomic constraints at every prediction level have not yet been systematically evaluated for diatom classification. Existing deep-learning-based diatom systems remain predominantly flat [gonzalez2025, diatomnet2024], while existing hierarchical diatom methods rely on hand-crafted features or non-differentiable ensembles [dimitrovski2012, chaushevska2020], leaving a gap between flat CNN architectures and end-to-end hierarchical models with explicit structural constraints in this domain.
@@ F-S Flat Species 2-layer MLP Backbone only Single focal
Figure 2: Hierarchical diatom classification architecture (H-COFGS). The EfficientNet-B0 backbone extracts features from input images, which are fed to five cascaded classification heads (Class, Order, Family, Genus, Species). Each head receives backbone features concatenated with probability distributions from all parent levels. Input diatom image adapted from [berezovska2016diatom], $C$ BY-SA 4.0.
TABLE II: Hierarchical masking example at the order level. Given predicted class Bacillariophyceae (pennate diatoms). Orders from class Bacillariophyceae receive mask value 1, while the ones from other classes are masked with -$\infty$, yielding zero probability after softmax. Softmax is computed as part of the focal loss during training.
Order Parent Class Logit Mask Masked Logit Probability
Naviculales Bacillariophyceae 4.2 1 4.2 0.730
Eunotiales Bacillariophyceae 2.8 1 2.8 0.180
Cymbellales Bacillariophyceae 1.5 1 1.5 0.049
Mastogloiales Bacillariophyceae 0.9 1 0.9 0.027
Cocconeidales Bacillariophyceae 0.3 1 0.3 0.015
Thalassiosirales Coscinodiscophyceae 3.1 0 -$\infty$ 0
Melosirales Coscinodiscophyceae 1.2 0 -$\infty$ 0
Fragilariales Fragilariophyceae 2.5 0 -$\infty$ 0
Tabellariales Fragilariophyceae 0.8 0 -$\infty$ 0
Sum of valid probabilities: 1.000
Transformation: Logit $\xrightarrow{\text{mask}}$ Masked Logit $\xrightarrow{\text{softmax}}$ Probability. Invalid orders ($\text{Mask}=0$) are set to -$\infty$, so $e^{-\infty}=0$ after softmax. Note: Thalassiosirales has a higher logit (3.1) than Cymbellales (1.5), but is masked out due to taxonomic constraints.
III Materials and Methods
III-A Dataset and Data Pipeline
III-A1 Data Cleaning and Taxonomy Construction
The initial raw dataset contains 4,881 microscope images extracted from the taxonomic reference work by Krammer and Lange-Bertalot [krammer2012] for research purposes. Taxonomic labels were manually verified and annotated by research staff and students at the Yantai Institute of Coastal Zone Research, Chinese Academy of Sciences, under the supervision of Dr. Yin-chu Wang. The dataset spans 4 classes, 15 orders, 28 families, 47 genera, and 684 species, following the taxonomic system of Round et al. [round1990].
Note: The image data is used solely for academic research and model training; no original images are published or distributed. All experimental results and visualizations in this paper use data derived from model predictions, not source images.
We apply systematic data cleaning to correct taxonomic inconsistencies: (1) merge the rare class Mediophyceae into Coscinodiscophyceae [round1990], (2) remove samples with incomplete annotations at class, order, or family levels, (3) exclude species with uncertain labels containing markers such as “sp.”, “cf.”, or “?”, and (4) resolve hierarchical conflicts where taxa map to multiple parents using majority voting. After cleaning, the dataset contains 3,522 samples across 3 classes, 13 orders, 23 families, 36 genera, and 608 species.
III-A2 Progressive Minimum Sample Filtering
To support progressive training and fair comparison, we apply a uniform threshold of 10 images per taxon at each taxonomic level, yielding four nested datasets: H-CO (3,506 samples, filters order), H-COF (3,495 samples, filters order and family), H-COFG (3,472 samples, filters order, family, and genus), and H-COFGS (1,456 samples, filters all four levels). The final filtered dataset, used by both hierarchical models H-COFGS and the flat baseline F-S, spans 3 classes, 9 orders, 13 families, 17 genera, and 82 species.
All datasets are randomly split into training (70%), validation (15%), and test (15%) sets, stratified by the deepest taxonomic level of each model (e.g., class for F-C, order for H-CO) with a fixed random seed (42) to ensure reproducibility. Input images are resized to $320\times 320$ pixels and normalized using ImageNet statistics. Data augmentation during training includes random horizontal flips, rotations within $\pm 15^{\circ}$, and color jittering [krizhevsky2012imagenet, shorten2019survey].
@@ III-B2 Hierarchical Models
Hierarchical models explicitly embed taxonomic relationships by predicting labels at multiple levels in a single forward pass through cascaded classification heads (Fig. 2). Our architecture incorporates two key mechanisms: cascaded probability propagation and hierarchical mask constraints.
Cascaded Probability Propagation. Prior hierarchical methods either use separate CNNs at each level [yan2015hdcnn] or parallel branches that predict levels independently without inter-level information flow [elhamod2022hierarchy]. In contrast, our cascaded architecture explicitly propagates probability distributions from parent levels to child levels. Each head at level $\ell$ receives concatenated features from the backbone and all parent-level probability distributions (Equation 1):
$\mathbf{x}_{\ell}=[\mathbf{f}_{\text{backbone}};\mathbf{p}_{\text{class}};\mathbf{p}_{\text{order}};\ldots;\mathbf{p}_{\ell-1}].$ (1)
This feature fusion provides a hierarchical context to each classification head, strengthening taxonomic consistency across levels. As illustrated in Fig. 2, head architectures scale with task complexity: the class head uses a two-layer MLP, intermediate levels (order, family) use three-layer MLPs, and fine-grained levels (genus, species) use four-layer architectures (detailed in Table I).
Hierarchical Mask Constraints. Another key design is the masking mechanism that enforces taxonomic consistency. At each level $\ell$, the model produces raw logits $\mathbf{z}_{\ell}$ transformed via masked softmax with a binary mask $\mathbf{m}_{\ell}\in\{0,1\}^{n_{\ell}}$:
@@
where $\mathbf{m}\ell$ is constructed by querying the taxonomy tree based on the parent label. Table II illustrates this mechanism. Table II illustrates this mechanism with a concrete example from our taxonomy.
III-C Training Strategy
III-C1 Multi-Level Loss Function
For hierarchical models, each taxonomic level receives a separate focal loss [lin2017focal] to address label imbalance, and the total loss is a weighted sum:
@@ III-C1 Multi-Level Loss Function
where weights increase progressively from coarse to fine levels (class: 0.8, order: 0.9, family: 1.0, genus: 1.2, species: 1.5) to emphasize harder, fine-grained predictions. Focal loss with $\alpha=0.25$ and $\gamma=2.0$ addresses label imbalance [lin2017focal, buda2018systematic].
III-C2 Training and Validation
All models are trained using the AdamW optimizer [loshchilov2019adamw] with an initial learning rate of $5\times 10^{-4}$, weight decay $1\times 10^{-4}$, a ReduceLROnPlateau scheduler (decay factor 0.5, patience 5 epochs), batch size 32, a maximum of 80 epochs, and early stopping with a patience of 15 epochs based on validation performance (deepest-level weighted F1 score).
Bidirectional Information Flow. Standard deep learning training typically follows four steps: (1) a forward pass that maps inputs to logits, (2) loss computation based on logits and ground-truth labels, (3) a backward pass that propagates gradients through the network, and (4) parameter updates performed by an optimizer. In our hierarchical setting, the forward pass is specialized: logits are computed sequentially from coarse to fine levels, and each head receives a feature vector obtained by concatenating the shared backbone representation with all previous levels’ probability distributions (Equation 1), so that hierarchical information flows explicitly from parent to child. During loss computation, we further exploit the taxonomy by applying ground-truth-derived hierarchical masks to the logits at each level and then feeding these masked logits into a weighted focal loss (Equation 3), effectively restricting every level’s candidate space to valid children of its parent and stabilizing training. The subsequent backward pass aggregates gradients from all levels and propagates them through the shared backbone, yielding bidirectional information flow: bottom-up via gradient backpropagation and top-down via the hierarchical constraints imposed during loss computation.
@@ III-C2 Training and Validation
We adopt a progressive training strategy across taxonomic depths. First, the flat class-only baseline F-C is trained from ImageNet initialization on the full cleaned dataset (3522 samples). Next, H-CO initializes its backbone from F-C and is trained on the class–order dataset H-CO (3506 samples), which is filtered to satisfy the minimum-sample threshold at the order level. Similarly, H-COF initializes its backbone from H-CO and trains on the H-COF dataset (3495 samples), H-COFG initializes from H-COF and trains on H-COFG (3472 samples), and H-COFGS initializes from H-COFG and trains on H-COFGS (1456 samples), each using its own filtered dataset that enforces the minimum-sample threshold at its deepest taxonomic level.
III-C3 Inference
During test inference, flat models predict a single taxonomic level directly, with labels at other levels obtained by querying the taxonomy tree. Hierarchical models use the same greedy hierarchical prediction as validation: logits are computed sequentially (Equation 1), probability distributions are computed via softmax, and we greedily select the category with the highest probability (argmax) at each level, using the predicted index from the previous level to mask the current level’s probability distribution via Equation 2, ensuring biological consistency: